OpenStudy (anonymous):

Find the equation of the line tangent to y = –3x2 + x – 5 at x = –1.

OpenStudy (welshfella):

first find the derivative dy/dx

OpenStudy (welshfella):

can you do that?

OpenStudy (anonymous):

Would it be -6x+1?

OpenStudy (welshfella):

yes so plug in x = -1 at this will give you the slope of the tangent

OpenStudy (welshfella):

* and this

OpenStudy (anonymous):

So the slope of the tangent line would be 7 right?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

now you need to use the general form y - y1 = m(x - x1) where m = 7 and (x1,y1) is a point on the line use x1 = -1 and find y1 by substituting in the equation of the curve.

OpenStudy (anonymous):

Would it be 2?

OpenStudy (anonymous):

Would 7x-2 be my answer?

OpenStudy (welshfella):

for y1? no y1 = -3(-1)^2 - 1 - 5

OpenStudy (anonymous):

Oh you get -9

OpenStudy (welshfella):

y -(-9) = 7(x -(-1)) y + 9 = 7x + 7 y = 7x - 2

OpenStudy (anonymous):

Okay thank you!

OpenStudy (welshfella):

yw